# Colorful Rings of Partitions

### v1

@@ 8. More personal remarks
The independent research done in high-school had led to the theorem mentioned above and others like the Pentagonal number theorem (which I however could not prove then). The topic of partitions came up during the winter 2023/2024 in the context of geometry, when looking again at the discrete Sard theorem. That theorem has classically has been developed by Morse and Sard [47, 54]. It is pretty remarkable that one can do a similar result within finite mathematics: first define carefully what a $d$-manifold is (a finite simple graph for which every unit sphere is a $(d-1)i$-sphere, where a $d$-sphere is a $d$-manifold that becomes contractible if a vertex is taken away. In 2015, I noticed that one can define hypersurfaces in such a manifold by looking at all simplices on which the function changes sign [34]. The only requirement is that the function $f$ is locally injective (a coloring). During Fall 2023 this was then generalized to higher dimensions [41]
@@ 9. Rings of Partitions
This motivates to investigate the growth rate of partition primes, similarly as we are interested in the growth rate of rational primes. The ring of partitions allows also to do cryptology as we can produce finite rings from the ring of partitions. Given a rational integer $N\in\mathbb{N}$ we can for all $n\in\mathbb{Z}_{N}$ define partitions $n=(n_{1},\dots,n_{k})$ of $n$, where $n_{i}\in\mathbb{Z}_{N}$ and $\sum_{i}n_{i}=n$. Now, if $k>n$, we can reduce the value of the $n$ smallest entries by $1$.This is a finite commutative ring with zero element 00 and one element $(1)$.
9.3.
Similarly than finite rings in number fields it quickly gets large. Now, schemes like Diffie-Hellman key exchange or RSA public key systems could be tried in this modular ring of partition. Also cryptological algorithms to factor integers could be used [51]. A simple example is the Pollard $\rho$ algorithm. It requires to build in fast ways to form $a^{n}$ if $a\in\mathcal{P}$ and $n\in\mathbb{Z}$ and to build the greatest common divisor ${\rm gcd}(n,m)$ of two elements $n,m$ in $\mathcal{P}$.

### v2

@@ 8. More personal remarks
The independent research done in high-school had led to the theorem mentioned above and others like the Pentagonal number theorem (which I however could not prove then). The topic of partitions came up during the winter 2023/2024 in the context of geometry, when looking again at the discrete Sard theorem. That theorem has classically has been developed by Morse and Sard [47, 54]. It is pretty remarkable that one can do a similar result within finite mathematics: first define carefully what a $d$-manifold is (a finite simple graph for which every unit sphere is a $(d-1)i$-sphere, where a $d$-sphere is a $d$-manifold that becomes contractible if a vertex is taken away. In 2015, I noticed that one can define hyper-surfaces in such a manifold by looking at all simplices on which the function changes sign [34]. The only requirement is that the function $f$ is locally injective (a coloring). During Fall 2023 this was then generalized to higher dimensions [41]
@@ 9. Rings of Partitions
This motivates to investigate the growth rate of partition primes, similarly as we are interested in the growth rate of rational primes. Unexplored is whether one can use partitions for cryptology. Given a rational integer $N\in\mathbb{N}$, is there a finite commutative ring to work with? We have experimented with some set-up. One could try to use it for schemes like Diffie-Hellman key exchange or RSA public key systems in such a modular ring of partition. In any finite Abelian group, there are cryptological algorithms to factor integers could be used [51]. A simple example is the Pollard $\rho$ algorithm. It only requires to build in fast ways to form $a^{n}$ if $a\in\mathcal{P}$ and $n\in\mathbb{Z}$ and to build the greatest common divisor ${\rm gcd}(n,m)$ of two elements $n,m$ in $\mathcal{P}$.